3X+7Y=-6
X-2Y=11
I AM LOOKING FOR HELP WITH SUBSTITUTION

Answer by checkley75(3666)

X-2Y=11
X=2Y+11 NOW SUBSTITUTE (2Y+11) FOR X IN THE OTHER EQUATION
3(2Y+11)+7Y=-6
6Y+33+7Y=-6
13Y=-6-33
13Y=-39
Y=-39/13
Y=-3 ANSWER
PROOF
X-2*-3=11
X+6=11
X=11-6
X=5 ANSWER
PROOF
3*5+7*-3=-6
15-21=-6
-6=-6